OpenStudy (anonymous):

What is the radius of a circle if the area is 113.04 inches squared?

OpenStudy (anonymous):

A = pi * r^2 113.04 = 3.14 * r^2 divide both sides by 3.14 36 = r^2 sqrt both sides sqrt(36) = sqrt(r^2) 6 = r